OpenStudy (anonymous):

(-2)^-1 I always suck at these

jigglypuff314 (jigglypuff314):

for example x^(-1) = 1/x

jigglypuff314 (jigglypuff314):

so (-2)^(-1) = 1/(-2)

OpenStudy (anonymous):

^-1 means inverse so the inverse of a whole number is a fraction so that -2^-1 means 1/-2